OpenStudy (rissaunnie):

i need three ordered pairs -x-2y=-10

OpenStudy (pawanyadav):

Why not you try with different values of x

OpenStudy (pawanyadav):

Assume any value of x and find corresponding value of Y by putting in the equation

OpenStudy (daggiemib):

-x-2y=-10 +x +x -2y=x-10 divide all by -2 y=-1/2x +5 graph it and you'll find ordered pairs

OpenStudy (pawanyadav):

Let's check for x=0 You get -2y=-10 So. y=5 Then one order pair is (0,5) Similarly find other

OpenStudy (rissaunnie):

like -10,0?

OpenStudy (daggiemib):

like y= -1/2(you choose a number) +5 then you can solve for y

OpenStudy (rissaunnie):

icant find a 3rd

OpenStudy (daggiemib):

try using zero?

OpenStudy (pawanyadav):

Put x=2,4,6. anyone you want

OpenStudy (daggiemib):

y=-1/2(2)+5 y=-1/2(4) +5 y=-1/2(6) +5